# Measuring Faithfulness of Chains of Thought by Unlearning Reasoning Steps

### v1

@@ Appendix D Add-mistake Implementation
We use gpt-4o-mini-2024-07-18 for inserting mistakes into each CoT step generated by our models. We then insert the modified CoT step containing the mistake in place of the original step and prompt each model to generate an answer based on the perturbed CoT, as per the prompt in Appendix B.
@@
Appendix E LLM-as-a-judge Setup
@@
Appendix F Additional Insights
@@
As mentioned earlier §6.2, we have found that efficacy correlates well with faithfulness. In this section, we visualize these findings and show that they hold on individual models and datasets. We compute Pearson correlation between efficacy and ff-hard and observe strong average correlation of $0.933$ with $p<0.0001$. We visualize the scatter plot of efficacy and faithfulness, measured as averages over all data points for each LR selection run (§C.1) in Figure 6. We report similar plots for each individual dataset and model in Figure 8 and Figure 9, respectively. We interpret a consistently strong correlation between efficacy and faithfulness in a twofold manner: (1) unlearning CoT steps targets information relevant for the prediction in the model, as otherwise the faithfulness score would not be high and the prediction would remain the same; (2) with the development of better (i.e. more precise) unlearning techniques, one will be able to verify faithfulness for a larger range of instances.
@@
Appendix G User Study
@@
Appendix H Hardware, Duration and Costs
@@
Generating data for the Add-mistake baseline (§D) was slightly more time consuming due to the few-shot prompting setup. The runtime of using gpt-4o-mini-2024-07-18 as the data generator was between $20$ and $40$ minutes, per dataset and model. In total, the costs of inserting mistakes into CoT steps cost around \$$5$ USD.
Appendix I Potential Risks
Our method aims to detect faithful reasoning steps in generated CoTs of LMs by unlearning information within those reasoning steps. We foresee two potential risks of our approach. Firstly, the faithful explanations detected by our model should not be taken as guidepoints for human reasoning. As our user study has shown (§6.3, §G), reasoning steps that are faithful to models are usually not plausible to humans, and should be used carefully in high-stakes scenarios. Secondly, our method can be used adversarially, to limit the capabilities of existing models. Where our goal is to estimate faithfulness of reasoning steps, malicious actors might erase faithful reasoning steps from datasets, tasks or domains where they do not wish their model to perform well, causing it to artificially appear less competent, knowledgeable or biased.

### v2

@@ Appendix F Add-mistake Implementation
We use gpt-4o-mini-2024-07-18 for inserting mistakes into each CoT step generated by our models. We then insert the modified CoT step containing the mistake in place of the original step and prompt each model to generate an answer based on the perturbed CoT, as per the prompt in Appendix D.
@@
Appendix G LLM-as-a-judge Setup
@@
Appendix H Additional Insights
@@
As mentioned earlier §6.2, we have found that efficacy correlates well with faithfulness. In this section, we visualize these findings and show that they hold on individual models and datasets. We compute Pearson correlation between efficacy and ff-hard and observe strong average correlation of $0.933$ with $p<0.0001$. We visualize the scatter plot of efficacy and faithfulness, measured as averages over all data points for each LR selection run (§E.1) in Figure 6. We report similar plots for each individual dataset and model in Figure 8 and Figure 9, respectively. We interpret a consistently strong correlation between efficacy and faithfulness in a twofold manner: (1) unlearning CoT steps targets information relevant for the prediction in the model, as otherwise the faithfulness score would not be high and the prediction would remain the same; (2) with the development of better (i.e. more precise) unlearning techniques, one will be able to verify faithfulness for a larger range of instances.
@@
Appendix I User Study
@@
Appendix J Hardware, Duration and Costs
@@
Generating data for the Add-mistake baseline (§F) was slightly more time consuming due to the few-shot prompting setup. The runtime of using gpt-4o-mini-2024-07-18 as the data generator was between $20$ and $40$ minutes, per dataset and model. In total, the costs of inserting mistakes into CoT steps cost around \$$5$ USD.
Appendix K Potential Risks
Our method aims to detect faithful reasoning steps in generated CoTs of LMs by unlearning information within those reasoning steps. We foresee two potential risks of our approach. Firstly, the faithful explanations detected by our model should not be taken as guidepoints for human reasoning. As our user study has shown (§6.4, §I), reasoning steps that are faithful to models are usually not plausible to humans, and should be used carefully in high-stakes scenarios. Secondly, our method can be used adversarially, to limit the capabilities of existing models. Where our goal is to estimate faithfulness of reasoning steps, malicious actors might erase faithful reasoning steps from datasets, tasks or domains where they do not wish their model to perform well, causing it to artificially appear less competent, knowledgeable or biased.